### v1

# MapGPT: Map-Guided Prompting for Unified Vision-and-Language Navigation

Abstract
Embodied agents equipped with GPT as their brain have exhibited extraordinary thinking and decision-making abilities across various tasks. However, existing zero-shot agents for vision-and-language navigation (VLN) only prompt the GPT to handle excessive environmental information and select potential locations within localized environments, without constructing an effective “global-view” (e.g., a commonly-used map) for the agent to understand the overall environment. In this work, we present a novel map-guided GPT-based path-planning agent, dubbed MapGPT, for the zero-shot VLN task. Specifically, we convert a topological map constructed online into prompts to encourage map-guided global exploration, and require the agent to explicitly output and update multi-step path planning to avoid getting stuck in local exploration. Extensive experiments demonstrate that our MapGPT is effective, achieving impressive performance on both the R2R and REVERIE datasets (38.8% and 28.4% success rate, respectively) and showcasing the newly emerged global thinking and path planning capabilities of the GPT model. Unlike previous VLN agents, which require separate parameters fine-tuning or specific prompt design to accommodate various instruction styles across different datasets, our MapGPT is more unified as it can adapt to different instruction styles seamlessly, which is the first of its kind in this field.
1 Introduction
Figure 1: A comparison of the thinking process of the GPT agent without and with topological maps. Given only a local action space, the agent may explore aimlessly, especially when navigation errors have already occurred. Incorporating topological maps enables the agent to understand spatial structures and engage in global exploration and path planning.
Large language models (LLMs) [27, 8, 41, 42, 14, 6] have demonstrated strong performance in various domains. As the most powerful LLMs, the GPT series models [8, 27, 28, 29] can even serve as the brain of embodied agents [21, 43, 26, 1], enabling them to engage in explicit thinking and decision-making process. Moreover, these GPT-based agents are typically zero-shot or few-shot, eliminating the burdensome tasks of data annotation and model training, and they also demonstrate remarkable generalization abilities.
Recently, LLMs have also been adopted in vision-and-language navigation (VLN), where agents are given human instructions that require them to visually perceive and navigate an indoor environment. Previous learning-based VLN methods [5, 16, 31, 12, 33, 45] relied on the training on large-scale domain-specific data with expert instruction annotations to execute navigation tasks. To address their reliance on training data and poor generalization, some GPT-based zero-shot agents [48, 24] with explainable decision-making abilities have been proposed to tackle the R2R dataset [5]. These methods focused on translating visual observations into textual prompts and required GPT-4 to act as an agent to select the correct position or direction.
However, such zero-shot VLN agents face several challenges. Firstly, transforming all visual observations into text may incorporate excessive useless and distractive information, thereby affecting the performance of the agent. Secondly, to enhance performance, these methods employ multiple additional experts to handle various subtasks, such as summarizing the history and fusing repeated expert predictions, which significantly increases the consumption of time and computational resources. Most importantly, these agents make decisions based solely on the observations of the local environment. This often makes them stuck in local exploration. As shown in Figure 1 (left), given only the local action space, when an agent realizes that it has engaged in an erroneous exploration, it can only continue to explore surrounding environment aimlessly.
We believe that the fundamental reason behind the struggling performance of these zero-shot agents lies in the absence of “global-view” to assist the agent in understanding the spatial structure of the real world, thereby hindering an efficient global exploration. We therefore introduce a simple yet efficient prompt system with only one navigation expert that prioritizes the processing of key node information (i.e., navigable viewpoints) in the environment. We further propose a map-guided prompting method for the GPT model to construct a zero-shot agent, named MapGPT, which aims to provide a unified solution for the VLN task. Specifically, we convert an online constructed topological map into textual prompts and feed them to the GPT model to assist in building and understanding the real-world navigation environment. Leveraging this map-guided prompt, our agent can engage in a more rational global exploration, as shown in Figure 1 (right). Interestingly, by applying the map-guided prompt, we have also uncovered a previously unknown capability of the GPT model, namely multi-step path planning. We require the agent to explicitly perform multi-step path planning and feed the planning as a reference input to the subsequent step, allowing iterative updates to the planning. As a result, the agent can perform explainable global planning to determine the exploration goals for several upcoming steps, or discover an appropriate backtracking path to a prior node for re-exploration if the ongoing exploration encounters obstacles.
We demonstrate impressive performance of MapGPT on the R2R [5] and REVERIE [32] datasets with different styles of instructions. Benefiting from the map-guided prompting and planning, our method outperforms all the existing zero-shot agents on the R2R significantly (around 4% improvement in success rate). In the unexplored zero-shot setting of the REVERIE, our agent exhibits enhanced competitiveness (28.4% success rate), surpassing even some learning-based methods trained on the REVERIE and only being inferior to certain approaches that incorporate additional information, such as precise GPS coordinates.
Benefiting from the flexible single-expert system, as well as the general map-guided prompts for navigation tasks, our MapGPT is more unified. With only some necessary prompt modifications, it can adapt to varying styles of instructions, without the need for extra decomposition of step-by-step instructions (not applicable to high-level instructions) as required by zero-shot DiscussNav [24], or separate parameter fine-tuning on individual datasets as in previous learning-based methods. Therefore, our MapGPT is the first unified approach in the VLN field, allowing effortless adaptation to different downstream tasks in this domain. Besides, MapGPT is applicable to multiple LLMs (e.g., GPT-4 and GPT-3.5) and we also develop the first VLN agent based on the recently released GPT-4V.
Our contributions can be summarized as follows.
  • We propose a novel map-guided path-planning agent, MapGPT, which converts a map constructed online into prompts for GPT, encouraging global exploration.
  • MapGPT achieves impressive performance on the R2R and REVERIE datasets, surpassing existing zero-shot agents. Besides, it is the first unified agent capable of adapting to varying instruction styles effortlessly.
  • By utilizing our novel map-guided prompting method, we uncover a previously unknown ability of the GPT model, namely multi-step path-planning in the real world.
2 Related Work
Vision-and-Language Navigation (VLN). As a representative multi-modal embodied AI task, VLN requires an agent to combine human instructions and visual observations to navigate and locate targets in real-world scenes. Previous learning-based approaches [44, 25, 15, 31] proposed various model architectures and trained their models on domain-specific datasets. Besides, pretrained models [20, 11, 12, 33, 18, 2, 34, 45, 30] have also been widely applied to this task to produce better multi-modal representations. Recently, to address the reliance on domain-specific data and explore a general agent in diverse scenarios, some zero-shot agents based on GPT [48, 24] have recently been proposed. However, they suffer from several limitations. For example, NavGPT [48] has limited performance and relies on a complex prompt system to process various environmental information. DiscussNav [24] introduces a sequential multi-experts system to discuss environmental information and fuses five repeated predictions to improve performance. Both of these two methods employ multiple experts with different roles to handle sub-tasks such as summarizing observations and decomposing instructions. These designs limit the agent’s capability to only address step-by-step instructions in the R2R dataset [5], and have not been validated on other styles of instructions (e.g., REVERIE [32]). Besides, these agents are limited to local exploration as they can only reason and make decisions within adjacent navigable points. In this paper, we propose a unified and efficient map-guided prompting for global exploaration and multi-step path planning, achieving impressive performance on both R2R and REVERIE.
Large Language Models (LLMs). LLMs [27, 28, 29, 46, 8, 13, 41, 42, 6] have demonstrated remarkable capabilities in multiple domains. Recently, LLM-based agents [21, 43, 26, 1, 47, 37, 30, 7, 36] have also attracted significant interest of the AI community. For example, VoxPoser [21] utilizes LLM and vision-language models to extract affordances and constraints, which enables motion planners to generate trajectories for manipulation tasks. SayCan [1] integrates language models and affordance functions to generate practical strategies for accomplishing the household task. LangNav [30] employs LLMs for navigation, but it merely utilizes GPT-4 [27] to synthesize some data and performs fine-tuning using Llama2 [42] as the backbone, rather than directly employing LLM as a zero-shot agent. In fact, the application of LLM-based zero-shot agents in navigation tasks is still limited, and how to prompt LLM to activate some crucial abilities required by navigation task have yet to be explored. In this paper, we propose a map-guided prompting method and uncover a previously unknown multi-step path-planning ability of the GPT models.
Maps for Navigation. Maps used for navigation tasks can be primarily categorized into two types, i.e., metric maps and topological maps. Employing SLAM [17] for constructing metric maps [9, 40] is widely used in navigation. However, this type of approach requires a trade-off between map size and computational efficiency, which affects navigation performance. To address this limitation, graph-based topological maps [10, 12, 2] have been proposed for pre-exploring environment or enabling global exploration, such as backtracking to previously visited nodes. However, these methods are all designed for model learning. It remains unexplored how to construct a map prompt and leverage the powerful capabilities of LLMs for zero-shot reasoning and planning based on the map.
Figure 2: The framework of our proposed MapGPT. Our basic system consists of two types of prompts, namely task description and fundamental inputs for VLN. We further introduce a map-guided prompting method that converts an online-constructed topological map into prompts, activating the agent’s global exploration. Finally, we require the agent to explicitly update multi-step navigation planning for each step based on the previous one, to enhance navigation performance.
3 Method
In this section, we first introduce our newly designed prompt system for the VLN task (Section 3.1). We then describe the details of our novel map-guided prompting method to establish the agent’s understanding of the global spatial structure of the environment (Section 3.2), which encourages the agent to explicitly generate and update its multi-step planning (Section 3.3).
3.1 Unified Single Expert Prompt System
Designing appropriate prompts for LLMs to activate their various abilities is of critical importance. However, prior NavGPT [48] and DiscussNav [24] gathered excessive noisy environmental information from all the views, which may interfere with the GPT’s decision-making. Besides, they relied on complex multi-expert systems, where GPT [27] played different roles to achieve various functions, such as instruction parsing, summarizing observations and history, etc. This undoubtedly increases resource consumption and makes it challenging to achieve seamless collaboration among multiple experts with different backgrounds. While a single expert may have a comprehensive understanding of complex task background and is capable of handling various sub-tasks uniformly.
As a comparison, our designed single expert prompt system has several features. (1) Utilizing one single navigation expert, we eliminate the need for a separate design of additional historical summary expert and progress estimation expert based on GPT models. (2) It is unified and flexible. We do not contain any customized design specific to instruction style, and it is convenient to incorporate additional information, such as maps. (3) We prioritize the observations of navigable points, which are combined with the corresponding directions to form a continuous semantics in action space, facilitating comprehension and decision-making.
As shown in Figure 2, we adopt a modular design for our prompt system. At each step, we collect various fundamental inputs for the agent, including instruction $I$, history $H_{t}$, observation $O_{t}$, and action space $A_{t}$. The meaning of these inputs, as well as the requirements for output, are clearly pre-defined in the task description $D$. We utilize a prompt manager $PM$ to organize these prompt inputs which are then fed into the large language model $LLM$ to generate the current thought $T_{t}$ and select a specific action $a_{t,i}\in A_{t}$. The pipeline can be formulated as
$T_{t},a_{t,i}=LLM(PM(D,I,H_{t},O_{t},A_{t})).$ (1)
Details will be provided in the following subsections.
3.1.1 Task Description
As shown in Figure 2, our prompts for task description $D$ consist of three main parts. Firstly, we set the task background by telling GPT that “You are an embodied robot that navigates in the real world …”. Secondly, we provide some definitions regarding the inputs at each step. For example, “‘Action options’ are some actions that you can take at this step”. Thirdly, we propose some basic requirements regarding how the agent should accomplish this task, such as some format requirements and an emphasis of thoughtful consideration like “think about what to do next and why …”.
3.1.2 Fundamental Inputs
Instruction $I$
Since we aim to adapt both fine-grained and high-level instructions, we directly feed the raw instructions to GPT without any initial analysis or decomposition.
Visual Observation $O_{t}$
Our agent is equipped with an RGB camera to capture $M$ images of the environment. Unlike NavGPT [48] that gathers excessive noisy environmental information, we utilize only the views that contain navigable points, and translate these observations into scene and object descriptions to prompt GPT to perceive the real world. The only exception is the stopping condition in the REVERIE dataset, which requires finding certain objects that may not be located in navigable directions, such as a painting hanging on a wall. Therefore, we utilize the bounding boxes provided by the REVERIE to extract an additional object list of all the views, named “Surroundings”, to enable the agent to determine whether to stop. Since we do not specify the directions of these surrounding objects, they cannot be used for navigation and are only utilized for stopping the agent in the REVERIE dataset.
To translate the visual observations into text descriptions, we follow NavGPT [48] and utilize off-the-shelf vision models BLIP-2 [23] and Faster R-CNN [35, 4] to acquire scene description and object detection respectively. Finally, we formulate these observations that have been converted into text as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point and is formulated using the prompt template:
“<{scene}>, which also includes <{objects}>”.
Action Space $A_{t}$
NavGPT allows the agent to directly select a viewpoint and DiscussNav enables the model to predict a direction. However, we incorporate both directional phrases (e.g., turn left to) and corresponding observation of this direction into the action space for GPT to select from, making the decision-making process more intuitive.
Specifically, we take as an input the action space $A_{t}=\{a_{t,i}\}_{i=0}^{N}$ for GPT model. $N$ is also the number of navigable points and we additionally define $a_{t,0}$ as “A. stop” so that the agent has $N+1$ options in total. Each $a_{t,i}$ in the remaining $N$ options is formulated using the template:
{label} {directional_phrase} {$o_{t,i}$.
We have defined six directional concepts, namely “go forward to”, “turn left to”, “turn right to”, “turn around to”, “go up to”, and “go down to”, according to the directions of navigable viewpoints. For instance, $a_{t,1}$ could be “B. turn around to <a room with blue walls>, which also includes <bed, curtain, picture>”. At each step $t$, the agent only needs to choose one option $a_{t,i}\in A_{t}$. For the output format, we require the agent to simply provide a single option label, such as ”Action: B.”, which is easy to parse.
History $H_{t}$
We directly record all previous actions $a_{0}\sim a_{t-1}$ for history. The following prompt template is utilized for appending the actions into $H_{t}$:
“step 0: {$a_{0}^{\ast}$}, step 1: {$a_{1}^{\ast}$}, …, step t-1: {$a_{t-1}^{\ast}$},
in which $t\geq 1$ and $a^{\ast}$ denotes the selected action $a$ but with the option label removed. The initial history is defined as $H_{0}$ = “The navigation has just begun, with no history”.
3.2 Map-Guided Prompting
For the VLN task, previous work [49, 12, 3] has demonstrated the effectiveness of online constructed maps for navigation. However, how to construct the maps and transform them into a certain form to prompt LLMs has not been investigated in this domain. In fact, there is also limited research exploring the capability of LLM as an agent in understanding a real-world environment in the field of LLM. In this paper, we propose a novel map-guided prompting approach that encourages GPT agents to engage in global exploration of the environment, as shown in Figure 2(b).
Topological Mapping
In the setting of the VLN task, the agent has never explored the entire environment and must construct a map based on its own observations online. We store the map as a dynamically updated graph, following the graph-based method DUET [12]. At each step $t$, we record all observed nodes along the navigation trajectory and their connectivity into the graph $G_{t}=\{V_{t},E_{t}\}$, where $V_{t}=\{v_{t,i}\}_{i=1}^{K}$ are a series of $K$ observed nodes marked with the index $i$ in the order of observations. All the edges between these observed nodes are recorded in $E_{t}$. At any step $t$, given the current location, the simulator will provide several neighboring nodes that are currently navigable. These new nodes and edges will be utilized for updating the graph from $G_{t-1}$ to $G_{t}$.
3.2.1 Converting Map into Prompt
After we have obtained a topological connectivity graph representing the structure of the environment, the next step is to transform it into an appropriate prompt and add map annotations to form a complete map-guided prompt to help the agent understand the navigation environment. For each step $t$, we categorize all observed nodes in the environment into three types, namely (1) explored nodes $\{en_{j}\}_{j=0}^{t}$ (including starting node $en_{0}$ and current node $en_{t}$), (2) accessible nodes $\{an_{t}^{0},an_{t}^{1},...\}$, and (3) unexplored inaccessible nodes $\{un_{0},un_{1},...\}$.
Trajectory
As we have already marked each location during the navigation process, it follows logically that we can create a simplified trajectory prompt to help the agent understand its navigation path in the map and avoid repeated exploration as far as possible. For the explored nodes, we formulate our trajectory prompt using the template:
“Trajectory: Place {$en_{0}$} {$en_{1}$}{$en_{t}$},
where each $en_{j}$ corresponds to a node $v_{t,i}\in V_{t}$ stored in the order of observation. Thus, we consider $i$ as the ID of this place and fill it into the template to denote $en_{j}$.
Map Connectivity
We transform the connectivity between places into the form of prompts as well, aiding the agent in understanding the spatial structure of the scene. Since the connectivity can only be observed at explored nodes, we always start with “Place {$en_{t}$} is connected with …”. All IDs corresponding to the neighboring accessible nodes of these exploded nodes will be listed using the following template:
“Map:
Place {$en_{0}$} is connected with Places {$an_{0}^{0}$}, {$an_{0}^{1}$}, …
Place {$en_{1}$} is connected with Places {$an_{1}^{0}$}, {$an_{1}^{1}$}, …
Place {$en_{t}$} is connected with Places {$an_{t}^{0}$}, {$an_{t}^{1}$}, …”,
where all the nodes should be filled with the place ID. Note that this map connectivity does not need to be updated if the agent decides to backtrack and revisit some previously explored nodes.
Map Annotations
The final step involves adding descriptions to each node of this topological map, enabling the agent to refer to them for path planning. As we have already provided descriptions of currently accessible nodes in the action space, and the selected actions are also included in the history to form explored nodes, there is no need to repeat these descriptions. It is sufficient to simply add the location IDs in the action space at each step. Specifically, each $a_{t,i}$ in the action space $A_{t}$ is reformulated as
{label} {directional_phrase} Place {$an_{t}^{i}$}: {$o_{t,i}$.
The agent can therefore find the corresponding descriptions of explored nodes and accessible nodes in history $H_{t}$ and action space $A_{t}$, respectively. However, we still have some unexplored and currently inaccessible nodes that are important, especially when the agent encounters obstacles in exploration and needs to revisit previous nodes for re-exploration. It is essential to record descriptions of these inaccessible nodes, as they provide supplementary information to assist the agent in backtracking to the most suitable node. We record these descriptions using the following template:
“Supplementary Info:
Place {$un_{0}$}: <{scene_0}>
Place {$un_{1}$}: <{scene_1}> …”,
where we retain only the scene descriptions for the agent’s reference. This is in line with the idea of coarse-scale representation in DUET [12]. Additionally, if there are no such nodes present in the map currently, we will directly set the prompt as “Supplementary Info: Nothing yet”.
3.3 Multi-Step Path Planning
NavGPT and DiscussNav record the thinking process of the agent at each step. Despite employing another GPT expert for summarization, they still involve a significant amount of redundancy. This is also not consistent with human thinking, as we usually do not document every moment of our thoughts. Instead, we tend to document a work plan and update it as necessary.
Inspired by the above insight and benefiting from the utilization of maps, we directly devise a planning module that demands the agent to explicitly generate its own multi-step path planning and dynamically update it at each step. Concretely, the agent is required to combine the current observation and thought to update a new multi-step path planning. The entire process is iterative, where the planning output of the current step serves as the input to the next step, allowing the agent to refer to previous plans. Therefore, at step $t$, the agent should refer to the last planning $P_{t-1}$, which is formulated as a prompt:
“Previous Planning: {$P_{t-1}$},
where $P_{0}$ is set as “Navigation has just started, with no planning yet”.
We have simultaneously modified part of the “output requirements” in the task description $D$ as follows: “Based on your ‘Map’, ‘Previous Planning’ and current ‘Thought’, you also need to update your new multi-step path planning to ‘New Planning’ ”. Therefore, we can parse the new planning $P_{t}$ in each step of the LLM output. In summary, the proposed MapGPT that combines map $M_{t}$ and path planning $P_{t-1}$ can be defined as follows:
$\small T_{t},P_{t},a_{t}=LLM(PM(D,I,H_{t},O_{t},A_{t},M_{t},P_{t-1})).$ (2)
In addition to the widely-used thought and action on the previous agents, MapGPT outputs additional multi-step planning information. This can help the agent perform path planning more effectively and strive to execute or update it automatically. A detailed analysis of these capabilities is presented in Section 4.4.
[…]
NavGPT [48] GPT-3.5 2 w/ 8.02 26.4 16.7 13.0
MapGPT (Ours) GPT-3.5 1 w/o 8.48 29.6 19.4 11.6
DiscussNav [24] GPT-4 5 w/o 6.30 51.0 37.5 33.3
MapGPT (Ours) GPT-4 1 w/o 5.80 61.6 41.2 25.4
MapGPT (Ours) GPT-4V 1 w/o 5.84 56.0 44.9 35.9
Table 1: Results on 72 various scenes of the R2R dataset. “Exp#” refers to the number of GPT experts and “Dist” refers to the utilization of distance information.
4 Experiments
4.1 Experimental Settings
Datasets and Evaluation
We choose two datasets, R2R [5] and REVERIE [32], to validate our MapGPT since they have distinct instruction styles. R2R provides detailed step-by-step instructions while REVERIE only offers a high-level description of finding the target object, which usually requires more exploration in the environment. To unify the prompt system, we focus only on navigation performance, which involves finding the correct location or object to stop, while neglecting the object grounding sub-task in REVERIE. We therefore adopt several evaluation metrics for navigation, including Navigation Error (NE, the distance between agent’s final location and the target location), Success Rate (SR), Oracle Success Rate (OSR, SR given Oracle stop policy), and SR penalized by Path Length (SPL).
Implementation Details
Most of the experimental results of MapGPT are achieved using GPT-4 [27]. For visual observations, we follow NavGPT [48] and utilize BLIP-2 [23] to provide a caption for the observation, and employ Faster R-CNN [35] pretrained on the Visual Genome dataset [22, 4] to detect existing objects. To transfer our MapGPT to the REVERIE dataset, we only make some simple yet necessary modifications, which demonstrates that our MapGPT is unified in the VLN task. Additionally, as shown in Table 1, MapGPT can be applied to multiple LLMs and has built the first GPT-4V-based [28, 29] agent in the VLN field. More details can be found in Appendix A.
[…]
Training Seq2Seq [5] 7.81 28 21 -
Speaker Follower [16] 6.62 45 35 -
EnvDrop [39] 5.22 - 52 48
Pretraining PREVALENT [19] 4.71 - 58 53
Rec-VLN-BERT [20] 3.93 69 63 57
HAMT [11] 2.29 73 66 61
DUET [12] 3.31 81 72 60
BEVBert [3] 2.81 84 75 64
Zero-shot DUET (Init. LXMERT [38]) 9.74 7 1 0
NavGPT [48] 6.46 42 34 29
MapGPT (Ours) 6.29 57.6 38.8 25.8
Table 2: Results on the validation unseen set of the R2R dataset under the single-prediction setting. MapGPT outperforms two non-pretrained methods and the zero-shot NavGPT.
4.2 Experimental Results
Comparison on the R2R Dataset
As shown in Table 1, we follow the settings of NavGPT [48] and DiscussNav [24] to first evaluate the performance of our proposed MapGPT compared to these zero-shot agents on 72 scenes of the R2R dataset. Except for 11 scenes from the validation unseen set of R2R, we also adopt 61 scenes from the R2R training set for the zero-shot setting. We randomly sample one path for each scene, with three instructions per path, resulting in a total of 216 instructions. By following this setting, we can evaluate the performance of these zero-shot agents across various scenarios. In addition, DiscussNav incorporates a general prompt technique to enhance performance. Specifically, it repeats the generation process of GPT-4 for five times and then leverage an additional expert to fuse these output results. However, this undoubtedly increases both the time and API costs significantly. Therefore, to ensure a fair comparison, we only evaluate the performance of our MapGPT, and previous NavGPT and DiscussNav under the single-prediction setting.
Through prompting GPT-4 to perform the VLN task, our MapGPT with only one GPT expert, surpasses DiscussNav in three metrics, decreasing navigation error (NE) by 0.5, gaining a 3.7% improvement in SR, and even achieving a boost of over 10% in OSR. MapGPT (GPT-4 based) has limited performance on the SPL metric, which could be attributed to two factors. Firstly, converting visual observations into textual descriptions results in information loss, causing the agent to struggle in accurately finding the target to stop. Secondly, map-guided prompts encourage the agent to continue global exploration when encountering uncertain situations, which increases the length of the navigation path, thus impacting the SPL metric. On the other hand, the MapGPT based on GPT-4V avoids the issue of information loss and achieves a performance of 35.9% on SPL. We also feed our proposed map-guided prompts into GPT-3.5, and it outperforms NavGPT’s results based on GPT-3.5. These experiments can prove that our proposed map-guided prompting is unified across different LLMs.
[…]
Training Seq2Seq [5] w/o 8.07 4.20 2.84
RCM [44] w/o 14.2 9.29 6.97
SMNA [25] w/o 11.3 8.15 6.44
FAST-MATTN [32] w/o 28.2 14.4 7.19
Pretraining HAMT [11] w/o 35.4 31.6 29.6
DUET [12] w/ 50.0 45.8 35.3
Zero-shot MapGPT (Ours) w/o 42.6 28.4 14.5
Table 3: Comparisons on a randomly sampled subset from the validation unseen set of the REVERIE dataset. Note that we have retested the released HAMT and DUET on this same subset.
We further compare the navigation performance between the proposed MapGPT and previous NavGPT on a larger validation unseen set with 11 scenes and 783 trajectories. As shown in Table 2, due to data distribution shift, the success rate of MapGPT is slightly lower (38.8%) compared to the results on the 72 scenes. Compared to NavGPT (GPT-4 based), MapGPT (GPT-4 based) exhibits a noticeable reduction in NE, leading by 4.8% in SR, and a substantial 15.6% improvement in OSR. This may be attributed to our proposed map-guided prompt, which has raised the upper limit of agent navigation ability (OSR). However, due to the need for a fair comparison, we utilize the same vision models as NavGPT, which constrains the final success rate (SR) of the agent in identifying the targets.
Moreover, this standard evaluation set can also be utilized for comparison with previous learning-based agents. In Table 2, we categorize previous methods into two types, namely training on the R2R dataset only and fine-tuning a pretrained model on the training set of the R2R dataset. As a zero-shot approach without utilizing any VLN data for training, MapGPT has demonstrated significant advancements over two learning-based methods, showcasing its tremendous potential.
In summary, our proposed MapGPT is highly effective, as it can be applied simultaneously to three GPT models, achieving the best performance in a zero-shot setting. Besides, it only utilizes a single GPT navigation expert and does not include additional distance information.
[…]
$\checkmark$ $\times$ $\times$ $\times$ 6.62 50.0 27.8 15.3
$\checkmark$ $\checkmark$ $\times$ $\times$ 6.49 49.5 32.9 19.4
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 6.40 59.7 37.5 24.8
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 5.80 61.6 41.2 25.4
Table 4: The performance impacts of different visual information extraction methods, incorporating map-guided prompts, and performing explicit multi-step path planning.
Figure 3: An example on the REVERIE dataset showcases MapGPT’s various abilities, including map understanding and global exploration (highlighted in yellow), multi-step path planning and the updating mechanism (comparing green and blue highlights, MapGPT plans to explore places 8, 9, and 10 at step 2, and gradually updates the progress until step 4), as well as backtracking capability (place 10 is not accessible from 9, thus backtracking to place 1 first).
Comparison on the REVERIE Dataset
Benefiting from the flexible single-expert system and the utilization of a universal map-guided prompt, we can effortlessly transfer MapGPT to the REVERIE dataset, adapting to different instruction styles. The validation unseen set of the REVERIE dataset contains 3,521 instructions in total. Considering the API costs and easier comparison for future work, we randomly sample 500 instructions from the validation unseen set for the zero-shot setting of the REVERIE benchmark. Besides, we have retested the released HAMT and DUET models on these 500 instructions for a fair comparison.
As shown in Table 3, the zero-shot MapGPT (GPT-4 based) exhibits greater competitiveness on the REVERIE, significantly outperforming some training-only models across all metrics. Moreover, when compared to HAMT which benefits from the pretraining and fine-tuning process, MapGPT demonstrates a highly competitive performance as well, surpassing by 7.2% in the OSR metric, though experiencing a 3.2% drop in SR. MapGPT’s inferior performance compared to DUET could potentially be attributed to the utilization of additional precise GPS coordinate information in DUET, which is not adopted in our MapGPT.
4.3 Ablation Study
As shown in Table 4, we conduct an ablation study to explore the effectiveness of some crucial information in the prompt on 72 scenes of the R2R dataset. For basic prompt systems that do not employ map and planning, we discover that the extraction of visual information significantly impacts navigation experts. When we only include scene-level description, i.e., a caption, for the visual observation, our baseline can only obtain a success rate of 27.8%. However, with the inclusion of object detection, there is a significant improvement in success rate, reaching 32.9%, while the OSR remains unchanged. This could be attributed to the fact that the results of object detection assist the agent in better identifying the correct locations to stop, while providing limited assistance in the process of exploration.
Based on this baseline system, we first introduce map information, which significantly enhances the agent’s global exploration capability and improves the accuracy in OSR by over 10%. These performance gains also translated partially into the final task success rate, increasing it from 32.9% to 37.5%. Furthermore, we incorporate a planning iterative updating mechanism, enabling the agent to reference path planning from the last step. This mechanism does not significantly improve the OSR performance since the map already provides sufficient capability for global exploration. However, it influences the process of decision-making, raising the success rate from 37.5% to 41.2% ultimately.
4.4 Case Study
In Figure 3, we showcase an example from REVERIE that sufficiently demonstrates the various abilities brought to the agent benefiting from map-guided prompting and iterative updating of multi-step path planning. In step 2, the agent is located at Place 1. Benefiting from the map understanding, it astutely identifies that out of the five places connected to its current position, only three remain unexplored. Since all these three places are described as bedrooms, the agent makes a multi-step path planning to sequentially explore them to find the bathroom. By Step 4, the agent has successfully executed the previous planning and all planned places except Place 10 have been explored. However, as Place 10 is not accessible from Place 9, the agent demonstrates its global exploration based on the map and decides to backtrack to Place 1 to continue exploring Place 10. The path planning information is also transmitted to subsequent navigation steps through an iterative updating mechanism. The latest planning is “go back to Place 1 and then explore Place 10…”, while the previous planning at step 2 is “I will first go to Place 8, then Place 9, and finally Place 10…”. This avoids the interference caused by recording all thoughts of each step and ensuring accurate execution of multi-step path planning. If the agent encounters new situations, it can also autonomously decide to remake a new path planning.
5 Conclusion
In this paper, we propose a novel zero-shot agent, named MapGPT, for the VLN task. MapGPT utilizes map-guided prompts to activate the global exploration and path-planning capabilities of the LLMs, thus avoiding getting stuck in the local environment. Additionally, we require the agent to explicitly generate and iteratively update planning to further enhance navigation performance. Extensive experiments demonstrate the impressive performance of MapGPT on both the R2R and REVERIE datasets.
Appendices
Appendix A More Details
A.1 Prompts
We provide specific task description prompts that are directly fed into the system content of GPT [27, 8] API. As shown in Figure 4, our unified prompts consist of three parts, namely task background, input definitions, and output requirements.
We have achieved effortless adaptation between fine-grained R2R [5] instructions and high-level REVERIE [32] instructions with only a few intuitive and necessary modifications. These modifications are primarily utilized for ignoring the extensive interactive actions with objects in REVERIE, since our unified agent is designed to focus on the navigation task. The stopping conditions also differ, as R2R only requires the agent to stop at the destination, while REVERIE demands the agent to check the target object in its surroundings before stopping. Additionally, unlike R2R, REVERIE instructions contain a significant amount of details regarding landmarks and objects. Due to the limitations of the visual model’s capability, we are unable to fully acquire this information. Therefore, we modify the prompts to require the agent to ignore these details.
A.2 Implementation
We implement the MapGPT agent using the GPT [27, 28, 29] API and the “gpt-4-0613” version of the model is adopted for GPT-4-based MapGPT. The “gpt-3.5-turbo-0613” and “gpt-4-vision-preview” version of the models are employed for GPT-3.5-based and GPT-4V-based MapGPT, respectively. Additionally, we set the temperature parameter of GPT models to 0, and all experiments are performed with a single-prediction setting. Moreover, we employ the visual model BLIP-2 [23] to extract captions as scene descriptions, using the released “blip2-flan-t5-xl” model. It is worth noting that we utilize identical prompts for GPT-3.5 and GPT-4 models. As for GPT-4V, we make no additional modifications, only incorporate some descriptions of image modality and enable the agent to directly reference the provided images corresponding to places in the action space and history, eliminating the need for converting images into textual descriptions.
Appendix B Additional Experimental Results
[…]
MapGPT 5.80 61.6 41.2 25.4
   w/o Trajectory 6.34 59.7 38.4 23.8
   w/o Supplementary Info 6.52 59.3 37.5 24.2
   w/o Updating Mechanism 6.43 58.3 35.6 21.8
Table 5: Ablation on the R2R dataset about removing “Trajectory” or “Supplementary Info” in the map prompts, as well as the iterative updating mechanism of the multi-step planning.
In the main text, we have reported the results of removing all the map-related or planning-related prompts. Here, we conduct further ablation study on some design details of prompts related to maps and planning.
As shown in Table 5, We retain the core prompts of the map and planning, while separately removing the “Trajectory” or “Supplementary Info” from the map, as well as the iterative updating mechanism of multi-step planning. The trajectories assist the agent in understanding its navigation path within the map, while the supplemental information represents descriptions of nodes that are currently inaccessible. Removing these pieces of information affects the agent’s perception and decision-making regarding the environment, subsequently impacting its performance. Therefore, these prompt designs are necessary.
We also test requiring the agent to output only some multi-step planning at each step, which can be considered as another form of “thought”, instead of using an iterative updating mechanism to propagate planning information to the next step. It can be observed that this prompt design is harmful to navigation performance, as the agent engages in redundant long-term planning that is not utilized in the following steps. Therefore, it is necessary to iteratively update and guide the agent’s subsequent decisions based on multi-step planning.
Appendix C More Qualitative Examples
We provide additional successful and failure cases on the R2R and REVERIE datasets to qualitatively analyze the capabilities and limitations of our proposed MapGPT (GPT-4 based).
Figure 5 demonstrates a successful case on the R2R dataset. In step 6, after thoroughly exploring places 3 and 4 connected to place 1, the agent decides to backtrack to place 1 and subsequently explore currently inaccessible places 6 and 7. Ultimately, the agent successfully terminates at place 7 in step 10. Figure 6 shows another successful case on the REVERIE dataset. This example shares similar scenarios with the example in the main text, but it differs in terms of the starting point and the given instruction. Here, we present a more detailed navigation process. In step 4, the agent makes appropriate planning based on the map and sequentially explores places 11, 12, and 13 in search of a bathroom. In step 8, after exploring place 13, the agent discovers that place 14, which is connected to the current location, contains a toilet and sink, indicating a potential correct direction. Therefore, the agent continues exploring and finally stops at the correct destination.
As shown in Figure 7, we further summarize three typical types of failure cases. (a) The process of converting visual observations into textual descriptions may result in error or information loss. For example, the agent recognizes the room as a hallway, causing the agent to fail to stop at the correct location. (b) The agent may fail to follow the details in the instructions accurately. For instance, instead of walking straight into a bedroom in the eleven o’clock direction as instructed, it turns left in step 1 and enters another incorrect bedroom, and stops there. (c) The scenes are highly challenging, and the instructions may not provide many clues. Thus, the agent may fail to explore the correct direction in time. To address these limitations, one possible research direction for the future is to utilize increasingly powerful multimodal models and combine them with map-guided prompts for making multi-step planning directly, without the need to convert visual observations into textual descriptions.
Figure 4: Task description prompts for the R2R and REVERIE datasets. We make some simple yet necessary modifications to transfer the MapGPT from the R2R task to REVERIE. This work focuses on unified navigation, while instructions in REVERIE often require some interactive actions on objects. Therefore, we require the agent to ignore these actions, as well as various object details that are difficult for the vision models to extract adequately.
Figure 5: A successful example on the R2R dataset. We demonstrate some crucial steps that leverage map-guided global exploration and planning capabilities, ultimately resulting in successful navigation.
Figure 6: A successful example on the REVERIE dataset. We demonstrate some crucial steps that leverage map-guided global exploration and planning capabilities, ultimately resulting in successful navigation.
Figure 7: Three typical reasons for failure. (a) The error or loss of information due to the transformation from visual observation to textual description (the agent passes the correct location but fails to stop). (b) Stopping erroneously in similar locations. (c) Failure to timely backtrack and explore the correct direction if the scene is complex and the instruction does not provide much guidance (the agent has explored place 1 and plans to explore place 2, while the correct direction is 0$\rightarrow$3).